OpenStudy (steve816):

It has been 4 hours. My hope is slowly diminishing. Please MAYDAY! MAYDAY! Find the equation of a line containing the points S(-3,4) and T(6,-7) in slope-intercept, standard, and point-slope form.

jimthompson5910 (jim_thompson5910):

what slope did you get?

OpenStudy (steve816):

actually I think it's -9/11

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

no you had it right the first time

OpenStudy (steve816):

oh -11/9

OpenStudy (steve816):

So what do I do from here?

jimthompson5910 (jim_thompson5910):

\[\Large m = \frac{y_2-y_1}{x_2-x_1}\] \[\Large m = \frac{-7-4}{6-(-3)}\] \[\Large m = \frac{-7-4}{6+3}\] \[\Large m = \frac{-11}{9}\]

jimthompson5910 (jim_thompson5910):

now you use one of the points along with the slope

OpenStudy (steve816):

Yes thank you for that.

jimthompson5910 (jim_thompson5910):

and plug them into the point slope form \[\Large y - y_1 = m(x-x_1)\]

OpenStudy (steve816):

how do I find the y intercept?

jimthompson5910 (jim_thompson5910):

\[\Large y - y_1 = m(x-x_1)\] \[\Large y - y_1 = -\frac{11}{9}(x-x_1)\] \[\Large y - 4 = -\frac{11}{9}(x-(-3))\] \[\Large y - 4 = -\frac{11}{9}(x+3)\] In Step 2, I plugged in the slope In Step 3, I plugged in one of the given points. In this case, I used (-3,4)

jimthompson5910 (jim_thompson5910):

so hopefully you see how I got the point slope form equation

OpenStudy (steve816):

Wow, brilliant!

OpenStudy (steve816):

but how do I find the b value in slope-intercept form?

jimthompson5910 (jim_thompson5910):

now isolate y and convert the equation into y = mx+b form

jimthompson5910 (jim_thompson5910):

also you'll have to distribute the -11/9 through

OpenStudy (steve816):

Alright thanks!

jimthompson5910 (jim_thompson5910):

as for standard form, you have to convert to Ax + By = C form

jimthompson5910 (jim_thompson5910):

A,B,C are whole numbers

OpenStudy (steve816):

Yes I understand how to do this from there! Thanks so much, really appreciate it.

jimthompson5910 (jim_thompson5910):

you're welcome